The graph of the function f(x) is, and the the horizontal axis is x. Given the...

(1 point) The graph of the function f(x) is,

The graph of the function f(x) is, and the the horizontal axis is x.

Given the differential equation x?(t)=f(x(t))x?(t)=f(x(t)).
List the constant (or equilibrium) solutions to this differential equation in increasing order and indicate whether or not these equations are stable, semi-stable, or unstable. (The second line after the comma has a drop menu that asks for stable, semi-stable, or unstable.)

A.) ________ , ________

B.) ________ , ________

C.) ________ , ________

D.) ________ , ________
Answer #1

The equilibrium points are the points where the curve touches or crosses the horizontal axis.

So, the equilibrium points are x=-4, -1, 1 and 5/2

Now, f(x) is negative in (-inf, -4) and positive in (-4, -1) so, x=-4 is unstable.

f(x) is positive in (-4, -1) and positive in (-1, 1) so, x=-1 is semi-stable.

f(x) is positive in (-1, 1) and negative in (1, 5/2), so x=1 is stable.

f(x) is negative in (1, 5/2) and positive in (5/2, inf), so, x=5/2 is unstable.